# Leaking Queries On Secure Stream Processing Systems

### v1

@@ 4.1 Attack Overview
The attacker first instantiates the stream processing engine with a single operator running inside an enclave. This environment approximates the real, target environment, even though the attacker does not know which operator is running in the target application. As discussed in our system model (Section 3.1), the attacker knows the source code of the DSP system and the enclave code of the Job Manager and Task Manager. It then runs the system with one Job Manager enclave and one Task Manager enclave. It constructs a Job Graph consisting of one operator of its choice and sends it to the Job Manager enclave. The only difference between this environment and the real one is the Job Graph containing the specification of the operator (i.e., which operator and which parameters). In the attacker-controlled environment, the attacker can specify any operator with any parameters of its choice.
@@ 5.2.6 Query recovery
Query Operators Total Random Forest SVM XGBoost
@@ 7 Related Work
Apache Storm [4] is among the first distributed stream processing systems. It adopts the execution model in which the query is represented as a DAG of operators, and data flows from one operator to the next. Subsequent systems follow the same model, and adds more features to the original Storm design. Spark Streaming [7], Flink [6] and Samza [3] supports both batch and stream processing in a single framework, and includes many stateful operators such as joins and aggregates. These systems handle operator failures, which can affect query correctness, by replicating the operators, checkpointing [67], or a combination therefore [8]. Other works focus on improving the performance of DAG execution, by detecting bottleneck via backpressure [68], or by dynamically replicating the operators to adapt to workload changes [69, 70]. Our work demonstrates attacks against a generic DSP system, which can be applied to systems other than Flink.
@@
Designing an outsource database running on an untrusted cloud requires making trade-offs between security, functionality, and performance. Cryptographic solutions such as CryptDB [12] and Arx [71] protect data confidentiality using cryptographic techniques, namely using homomorphic and functional encryption schemes. However, they support limited computation over the data, and have been shown to leak data in practical settings [72]. Secure databases built on top of trusted hardware include Cipherbase [73] and TrustedDB [74] that use older hardware. More recent systems based on Intel SGX, such as ObliDB [75], Opaque [76], Obaldi [77], focus on obliviousness to prevent leakage through memory access pattern [78]. They achieve high performance and more functionalities than purely cryptographic solutions, but their security assumptions are stronger, i.e. the hardware is secure. Furthermore, they incur non-trivial overhead due to obliviousness.
@@ Appendix A Benchmark
In summary, our implementation of the NEXMark benchmark evaluates seven distinct stream processing operators (see Table VII). Notably, the Average operator appears in both Q4 and Q6, but with different windowing mechanisms: Q4 employs an Average operator over all data events, whereas Q6 uses a Partitioning, or Keying, average mechanism. Given that different aggregation strategies significantly affect operator behavior and processing characteristics, we treat the same operation with different aggregation mechanisms as separate operator types in our analysis.
@@ Appendix B Evaluation
TABLE VIII: Model Hyperparameters
@@
B.1 Hyperparameters
Table VIII lists the parameters of the models.

### v2

@@ 4.1 Attack Overview
The attacker first instantiates the stream processing engine with a single operator running inside an enclave. This environment approximates the real, target environment, even though the attacker does not know which operator is running in the target application. As discussed in our system model (Section 3.1), the attacker knows the source code of the DSP system and the enclave code of the Job Manager and Task Manager. It then runs the system with one Job Manager enclave and one Task Manager enclave. It constructs a Job Graph consisting of one operator of its choice and sends it to the Job Manager enclave. The only difference between this environment and the real one is the Job Graph containing the specification of the operator (i.e., which operator and which parameters). In the attacker-controlled environment, the attacker can specify any operator with any parameters of their choice.
@@ 5.2.6 Query recovery
Query Operators No. Samples Random Forest SVM XGBoost
@@ 7 Related Work
Apache Storm [4] is among the first distributed stream processing systems. It adopts the execution model in which the query is represented as a DAG of operators, and data flows from one operator to the next. Subsequent systems follow the same model and add more features to the original Storm design. Spark Streaming [7], Flink [6] and Samza [3] support both batch and stream processing in a single framework and include many stateful operators such as joins and aggregates. These systems handle operator failures, which can affect query correctness, by replicating the operators, checkpointing [67], or a combination thereof [8]. Other works focus on improving the performance of DAG execution by detecting bottlenecks via backpressure [68], or by dynamically replicating the operators to adapt to workload changes [69, 70]. Our work demonstrates attacks against a generic DSP system, which can be applied to systems other than Flink.
@@
Designing an outsourced database running on an untrusted cloud requires making trade-offs between security, functionality, and performance. Cryptographic solutions such as CryptDB [12] and Arx [71] protect data confidentiality using cryptographic techniques, namely, using homomorphic and functional encryption schemes. However, they support limited computation over the data, and have been shown to leak data in practical settings [72]. Secure databases built on top of trusted hardware include Cipherbase [73] and TrustedDB [74] that use older hardware. More recent systems based on Intel SGX, such as ObliDB [75], Opaque [76], Obaldi [77], focus on obliviousness to prevent leakage through memory access patterns [78]. They achieve high performance and more functionalities than purely cryptographic solutions, but their security assumptions are stronger, i.e. the hardware is secure. Furthermore, they incur non-trivial overhead due to obliviousness.
@@ Appendix A Benchmark
In summary, our implementation of the NEXMark benchmark evaluates seven distinct stream processing operators (see Table VIII). Notably, the Average operator appears in both Q4 and Q6, but with different windowing mechanisms: Q4 employs an Average operator over all data events, whereas Q6 uses a Partitioning, or Keying, average mechanism. Given that different aggregation strategies significantly affect operator behavior and processing characteristics, we treat the same operation with different aggregation mechanisms as separate operator types in our analysis.
@@ Appendix B Evaluation
B.1 Hyperparameters
Table VII lists the parameters of the models.
TABLE VII: Model Hyperparameters
@@ B.1 Hyperparameters
TABLE VIII: Summarization of deployed queries in NEXMark
No Summary Operators
Q1 Convert the value of each Bid from dollars to euros Map
Q2 Find Bids belong to specific Auctions, then return those Auctions and Bids’ price Filter, Map
Q3
Find People who are selling items of a specific category and their Auctions,
then return specific People’s information
Filter, Join, Map
Q4 Calculate the average prices of winning Bids of all Auctions Join, Max, Average
Q5 Count Auctions that have the most Bids Count
Q6 Calculate the average price per Person/seller Join, Filter, Max, Average (Partition)